Given the function
g(x) = 2 − x2,
evaluate
g(x + h) − g(x)
h
, h ≠ 0

Answer by Boreal(15235)

g(x+h)=2-(x+h)^2=2-x^2-2xh-h^2
g(x)=2-x^2
subtracting,
=-2xh-h^2
divide that by h and get -2x-h
as h approaches 0 get-2x
the derivative of 2-x^2 is -2x